OpenStudy (anonymous):

find to the nearest degree the angle whos measure is 2.5 radians

OpenStudy (anonymous):

\[\frac{\textrm{degrees}}{180}=\frac{\textrm{radians}}{\pi}\]

OpenStudy (anonymous):

thanks a bunch

OpenStudy (anonymous):

solve algebracially for x: 16 \[^{2x+3}\]=64\[^{x+2}\]